OpenStudy (anonymous):

Minimum and Maximum Problem: Can someone please help me?!? Where are the local max/min for the function e^x sin(x) between x=0 and x=2pi?

OpenStudy (anonymous):

take the derivative using the product rule set it equal to zero and solve when you set it equal to zero, factor out the \(e^x\) and since \(e^x\) is never zero, find the zeros of the remaining factor

OpenStudy (anonymous):

can you do it visually?

OpenStudy (anonymous):

no i would have to write it

OpenStudy (anonymous):

i don't really know what \(y=e^x\sin(x)\) looks like

OpenStudy (anonymous):

yea thats what I meant by writing it..

OpenStudy (anonymous):

\[f(x)=e^x\sin(x)\] \[f'(x)=e^x\sin(x)+e^x\cos(x)=e^x(\sin(x)+\cos(x))\]

OpenStudy (anonymous):

set \[\sin(x)+\cos(x)=0\] solve for \(x\)

OpenStudy (anonymous):

sine is the negative of cosine if \(x=\frac{3\pi}{4}\) or \(x=\frac{7\pi}{4}\)